• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Solvignon

    Log entry time 16:05:48 on March24,2011

    Entry number 353084

    keyword=Day Shift Summary

    08:00 Starting shift with production data on 20cm 4He at 4uA.

    08:45 Moved target to empty. MCC is going to send tuned beam before putting us in controlled access.

    14:23 Beam permit

    14:52 Tuned beam

    15:04 CW beam on BeO

    15:13 Moving to 20cm 4He

    15:22 Target in place. Asked for 4uA.

    15:31 LQ2 tripped

    15:56 LQ2 ready. Asked for beam back.

    16:00 Waiting for beam.



    Run summary:
    ------------
    * kinematics: Left Pe=3.6014 GeV/c, thetae=20.3deg
    Right Pp=1.3825 GeV/c, thetap=33.5deg
    I_BB = 518A


    run# 3021 4He Good 0.01954 C
    3022 4He Good 0.00444 C
    ---------
    Total= 0.02398 C



    * kinematics: Left Pe=3.6014 GeV/c, thetae=20.3deg
    Right Pp=1.30825 GeV/c, thetap=29.0deg
    I_BB = 518A

    run# 3025 4He Junk 0.0 C